OpenStudy (anonymous):

f(x)=x^3+3x^2-36x-108 Find x-intercepts of f(x)

OpenStudy (freckles):

to find x-intercepts set f(x)=0 and solve for x

OpenStudy (freckles):

\[x^3+3x^2-36x-108=0\]

OpenStudy (freckles):

You could try to factor by grouping and if that doesn't work we could try to guess a zero using some related theorem.

OpenStudy (anonymous):

I'm not sure what to do. I know to add 108 to each side though

OpenStudy (freckles):

Well I suggest factoring by grouping... we don't want to add 108 on both sides

OpenStudy (freckles):

look at x^3 and 3x^2 the first two terms what factor do they have in common?

OpenStudy (anonymous):

oh..i dont know how to factor by grouping

OpenStudy (anonymous):

x^2

OpenStudy (freckles):

So x^3+3x^2 can be written as x^2(x+3)

OpenStudy (freckles):

its look at the last two terms -36x-108

OpenStudy (freckles):

what do and -36 and -108 have in common

OpenStudy (anonymous):

-36

OpenStudy (freckles):

coolness so you can write -36x-108 as -36(x+3)

OpenStudy (freckles):

so we now have x^2(x+3)-36(x+3)

OpenStudy (freckles):

hmmm...

OpenStudy (freckles):

like we can now view this as if we have two terms

OpenStudy (freckles):

the first term is x^2(x+3) and the last term is -36(x+3)

OpenStudy (freckles):

I wonder if they have a common factor

OpenStudy (anonymous):

x+3

OpenStudy (freckles):

so you can factor (x+3) from both of those terms

OpenStudy (anonymous):

so..(x^2-36)(x+3) ?

OpenStudy (freckles):

bingo

OpenStudy (freckles):

So you can factor x^2-36 more but it really isn't necessary

OpenStudy (freckles):

we are ready to find our zeros

OpenStudy (freckles):

So we have (x^2-36)(x+3)=0 this happens when x^2-36=0 or when x+3=0

OpenStudy (freckles):

you need solve both of those equations

OpenStudy (anonymous):

-3, 6

OpenStudy (freckles):

one more solution

OpenStudy (anonymous):

-6

OpenStudy (freckles):

yep

OpenStudy (freckles):

x^2-36=0 x^2=36 when x=6 or when x=-6 since 6^2=6(6)=36 and (-6)^2=(-6)(-6)=36

OpenStudy (freckles):

So anyways you found your x-intercepts (-3,0) (6,0) (-6,0)

OpenStudy (anonymous):

thank you!!